# Reward Centering

### v1

@@ 3 Value-based Reward Centering
We draw inspiration from the average-reward formulation, where estimating the average reward in the off-policy setting is a pertinent problem. In particular, Wan et al. (2021) recently showed that using the temporal-difference (TD) error in average-reward RL (instead of the conventional error in (4)) leads to an unbiased estimate of the reward rate in the tabular off-policy setting. It turns out that this idea from the average-reward formulation is quite effective even in the discounted-reward formulation, which is the focus of this paper. We show that if the behavior policy takes all the actions that the target policy does (the exact distribution over actions may differ arbitrarily), then we get a good approximation of the average reward of the target policy using the TD error:
@@
The first column of Figure 3 shows plots for value-based centering in the on-policy problem from the previous section, where the target policy picks both actions with equal probability. Value-based centering (red) appears as good as simple centering (green) in terms of the rate of learning and asymptotic error. The other two columns show plots for two off-policy experiments with behavior policies $[b_{1}(\text{left}|\cdot),b_{1}(\text{right}|\cdot)]=[0.7,0.3]$, $[b_{2}(\text{left}|\cdot),b_{2}(\text{right}|\cdot)]=[0.3,0.7]$. The two different behavior policies are symmetric but result in different trends. Corresponding to $b_{1}$, we saw that value-based centering resulted in a lower RMSVE faster than simple centering for both values of $\gamma$, and the final error rate was roughly the same. As expected, the simple approach estimated the average reward incorrectly and hence the learned values were relatively larger than with value-based centering (but not as large as when there was no centering). The results with $b_{2}$ were more interesting. The RMSVE reduced rapidly at first with simple centering, then rose sharply, and then reduced again. This is because the average-reward estimate was initialized to zero and it converged to around $0.5$ (because $b_{2}$ skews the agent’s state distribution towards the more-rewarding right side). When the estimate passed the true value of $0.25$, the RMSVE was quite low, however, the estimate quickly climbed to $0.5$, resulting in the peak in RMSVE. Eventually the value estimates settled to values corresponding to an average-reward estimate of around $0.5$. In contrast, the average-reward estimate was much closer to the true value when using value-based centering, resulting in a smoother learning curve. The effects were amplified with the larger discount factor (bottom row).
Overall, we observed that reward centering can improve the rate of learning of discounted-reward prediction algorithms such as TD-learning, especially for large discount factors. While the simple way to center rewards is quite effective, value-based reward centering is better suited for general off-policy problems. We shall now consider reward centering within the control setting.
@@ 4 Case Study: Q-learning with Reward Centering
Experiments: We now present results of Q-learning with and without centering on a set of control problems with tabular, linear, and non-linear function approximation. The problems are primarily from CSuite (Zhao et al., 2022). The repository specifies each problem in detail; we provide high-level descriptions here. We start the assessment in a tabular problem and then proceed to problems that require function approximation.
@@
From the results so far, we observed that reward centering can improve the performance of tabular, linear, and non-linear variants of the Q-learning algorithm on various problems. The improvement in the rate of learning is larger for discount factors close to 1. Furthermore, there is an improvement in the robustness of the algorithms to shifts in the problems’ rewards. The parameter studies in this section indicate that the benefits of reward centering are quite robust to the choice of its parameter $\eta$. Appendix C contains additional learning curves and parameter studies that further reinforce the trends observed in this section.
@@ 5 Discussion, Limitations, and Future Work
The reward-centering idea is very general can be added to any discounted-reward RL algorithm for solving continuing problems. We focused on the reward-centered versions of TD-learning and Q-learning in this paper; we expect similar trends for all other algorithms that estimate values, like Sarsa (Rummery & Niranjan, 1994) or the family of actor-critic methods [^1] The advantage function or the baseline in policy-based methods have orthogonal benefits to reward centering. We discuss these in Appendix D. (Konda & Tsitsiklis, 1999; Schulman et al., 2016). Our preliminary experiments of adding reward centering to the PPO algorithm (Schulman et al., 2017) showed improvements on several Mujoco problems (Todorov et al., 2012) that we made continuing (see Appendix C). A comprehensive empirical investigation of reward centering added to other RL algorithms is an exciting direction of future work. In parallel, it is also pertinent to extend theoretical results from the tabular setting to that of function approximation, beginning with linear function approximation.
Reward centering is related to but different from simply learning a bias parameter. In theory, if a bias weight (corresponding to a bias feature of one) is equal to the state–action-independent offset $r(\pi)/(1-\gamma)$, then the rest of the function-approximation capacity can estimate the relative values. However, on testing this we found the learned bias weight was rarely close to true value of the offset. Moreover, it is desirable to separate $r(\pi)$ from the scaling effect of $1/(1-\gamma)$. Along these lines, Tsitsiklis and Van Roy (2002) suggested setting a linear approximator’s bias feature proportional to $1/(1-\gamma)$ and learning the corresponding bias weight. Reward centering takes this idea forward. Since we know what the corresponding bias weight is, we can have an explicit update rule for it (à la Sutton (1988b)), which can lead to faster learning. Besides, in the control problem, there is no need to estimate the offset within the standard discounted values, as Tsitsiklis and Van Roy propose doing. Instead, via reward centering, we can simply estimate the centered (relative) values.
Note that reward centering adds an additional element of non-stationarity to the learning problem because the average-reward estimate changes over time. As a result, the asymptotic variance of the updates is larger compared to when there is no centering (see, e.g., Figure 3). The increased variance of updates also ties in to how the average reward is estimated. Ideally, the average reward is accurately estimated as quickly as possible and does not change too much per step. The simple centering technique quickly estimates the average reward but is affected by the per-step stochasticity. Value-based centering additionally relies on changing values and hence can be relatively slower, but because it relies on the value estimates, per-step stochasticity in the problem dynamics may not affect it as much. [^2] Consider a fully deterministic tabular problem. The rewards are different at each step in general and so the simple centering technique would always involve non-zero errors. On the other hand, the per-step TD error would be zero in the value-based approach. A step-size adaptation technique would be pertinent for estimating the average reward: larger step sizes when the errors are consistently large, smaller step sizes otherwise. Such an adaptation should be possible throughout the lifetime of the agent, since it may encounter new parts of the world having different kinds of reward signals.
Speaking of different kinds of reward signals, reward centering should be combined with approaches that deal with different scales of rewards. Reward centering makes learning algorithms robust to shifts in the rewards. However, the relative values themselves may have large magnitude (e.g., if all the rewards in a problem are multiplied by a constant, then the relative values will be scaled by the same constant). It is thus pertinent to combine reward centering with reward scaling. The general idea of scaling quantities to a tractable range is quite common (e.g., van Hasselt et al., 2016; Pohlen et al., 2018). Building on this body of work, Schaul et al. (2021) recently proposed a lightweight trick that re-scales TD errors into an optimization-friendly range using problem-agnostic information in the agent’s data stream. With some care, their techniques can likely be extended from the episodic to the continuing case.
@@
One of the most intriguing follow-up directions based on reward-centering is a method that adapts the discount factor over time. Learning both the average reward and the centered discounted estimates allows the agent to rapidly estimate the discounted value function corresponding to any discount factor. Concretely, consider the agent has estimated the average reward $\bar{R}$ and the centered discounted value function $\tilde{{\mathbf{v}}}^{\gamma_{1}}$ to some level of accuracy. With just this information, the agent can form an estimate of the standard discounted value function corresponding to another discount factor $\gamma_{2}$ via $\bar{R}/(1-\gamma_{2})+\tilde{{\mathbf{v}}}^{\gamma_{1}}$. This is an estimate, of course, but it can be improved quickly with a few samples of experience—potentially with old experience from a buffer or a parameterized model. In contrast, with standard methods, it would take comparatively longer to raise the estimates to the new mean value and adapt the relative values. Hence, with reward centering, we can imagine efficient methods that adapt their discount factors over time: a low discount rate to learn quickly amidst a lot of uncertainty—like in the beginning of training—and when the world is more predictable, a higher discount rate to estimate the policy that maximizes the total reward obtained by the agent.
Reward centering is a simple idea with potentially large benefits. When combined with appropriate step-size adaptation and reward-scaling techniques, we think it will be a key enabler for agents to learn quickly and continually over their lifetimes.
@@ Acknowledgments
We would like to thank Huizhen Yu, Arsalan Sharifnassab, Khurram Javed, and other members of the RLAI lab for discussions that helped improve the quality and clarity of the paper. We are also grateful for the computing resources provided by the Digital Research Alliance of Canada.
@@ 14      ${\mathbf{x}}={\mathbf{x}}^{\prime}$
We recommend two small but useful optimizations to these general pseudocodes in Appendix C.
@@ Appendix D Connections to Related Approaches
Reward centering and the advantage function have orthogonal benefits. The advantage function benefits the actor by reducing the variance of the updates in the policy space (Sutton & Barto, 2018; Schulman et al., 2016). On the other hand, reward centering benefits the critic’s or baseline’s estimation by eliminating the need to estimate the large state-independent constant offset. Both the quantities involved in the advantage function—${a}_{\pi}^{\gamma}(s,a)={q}_{\pi}^{\gamma}(s,a)-{v}_{\pi}^{\gamma}(s)\ \forall s,a$—have the large state-independent offset $r(\pi)/(1-\gamma)$. The net effect of the offset is zero when they are subtracted. But the key point is that both the state- and action-value estimates include the large offset. Reward centering removes the need to estimate the large offset for both the state- and action-value function, which simplifies the critic-estimation problem. The actor update is left unchanged with reward centering because the advantage function itself remains unchanged: $\tilde{a}_{\pi}^{\gamma}(s,a)=\tilde{q}_{\pi}^{\gamma}(s,a)-\tilde{v}_{\pi}^{\gamma}(s)$, because $\tilde{q}_{\pi}^{\gamma}(s,a)=q_{\pi}^{\gamma}(s,a)-r(\pi)/(1-\gamma)$ and $\tilde{v}_{\pi}^{\gamma}(s)=v_{\pi}^{\gamma}(s)-r(\pi)/(1-\gamma)$. Hence, we expect reward centering to benefit all the algorithms that estimate values, which include all actor-critic methods that involve advantage estimation.
@@
Reward centering can be seen as reward shaping (Ng et al., 1999) with a constant state-independent potential function: $\Phi(s)=r(\pi)/(1-\gamma)\ \forall s$. Their Theorem 1 then reiterates that reward centering does not change the optimal policy of the problem. A possible drawback of reward shaping is that fully specifying the potential-based shaping function can be tricky, especially for problems with large state spaces. In the case of reward centering this is relatively easy: the potential function is constant across the entire state space, and we know how to learn the average reward reliably from data.

### v2

@@ 3 Value-based Reward Centering
We drew inspiration from the average-reward formulation of reinforcement learning, where estimating the average reward in the off-policy setting is a pertinent problem. In particular, Wan et al. (2021) recently showed that using the temporal-difference (TD) error (instead of the conventional error in (4)) leads to an unbiased estimate of the reward rate in the tabular off-policy setting. It turns out that this idea from the average-reward formulation is quite effective even in the discounted-reward formulation, which is the focus of this paper. We show that if the behavior policy takes all the actions that the target policy does (the exact distribution over actions may differ arbitrarily), then we get a good approximation of the average reward of the target policy using the TD error:
@@
The first column of Figure 3 shows plots for value-based centering in the on-policy problem from the previous section, where the target policy picks both actions with equal probability. Value-based centering (red) appears as good as simple centering (green) in terms of the rate of learning and asymptotic error. The other two columns show plots for two off-policy experiments with behavior policies $[b_{1}(\text{left}|\cdot),b_{1}(\text{right}|\cdot)]=[0.7,0.3]$, $[b_{2}(\text{left}|\cdot),b_{2}(\text{right}|\cdot)]=[0.3,0.7]$. The two different behavior policies are symmetric but resulted in different trends. Corresponding to $b_{1}$, we saw that value-based centering resulted in a lower RMSVE faster than simple centering for both values of $\gamma$, and the final error rate was roughly the same. As expected, the simple approach estimated the average reward incorrectly and hence the learned values were relatively larger than with value-based centering (but not as large as when there was no centering). The results with $b_{2}$ were more interesting. The RMSVE reduced rapidly at first with simple centering, then rose sharply, and then reduced again. This is because the average-reward estimate was initialized to zero and it converged to around $0.5$ (because $b_{2}$ skews the agent’s state distribution towards the more-rewarding right side). When the estimate passed the true value of $0.25$, the RMSVE was quite low, however, the estimate quickly climbed to $0.5$, resulting in the peak in RMSVE. Eventually the value estimates settled to values corresponding to an average-reward estimate of around $0.5$. In contrast, the average-reward estimate was much closer to the true value when using value-based centering, resulting in a smoother learning curve. The effects were amplified with the larger discount factor (bottom row).
Overall, we observed that reward centering can improve the rate of learning of discounted-reward prediction algorithms such as TD-learning, especially for large discount factors. While the simple way to center rewards is quite effective, value-based reward centering is better suited for general off-policy problems. Next, we consider reward centering within the control setting.
@@ 4 Case Study: Q-learning with Reward Centering
Experiments: We present results of Q-learning with and without centering on a set of control problems with tabular, linear, and non-linear function approximation (see Appendix A for the pseudocode). The problems are primarily from CSuite (Zhao et al., 2022). The repository specifies each problem in detail; we provide high-level descriptions here. We start the assessment in a tabular problem and then proceed to problems that require function approximation.
@@
Through these experiments, we observed that reward centering can improve the performance of tabular, linear, and non-linear variants of the Q-learning algorithm on various problems. The improvement in the rate of learning is larger for discount factors close to 1. Furthermore, there is an improvement in the robustness of the algorithms to shifts in the problems’ rewards. The parameter studies in this section indicate that the benefits of reward centering are quite robust to the choice of its parameter $\eta$. Appendix C contains additional learning curves and parameter studies that further reinforce the trends observed in this section.
@@ 5 Discussion, Limitations, and Future Work
Reward centering can improve the data efficiency and robustness of almost any algorithm for continuing reinforcement learning. Here we have shown improvements for algorithms that learn state-value functions and action-value functions, and for algorithms that are tabular or use linear or non-linear function approximation.111 As a further example, in preliminary experiments in Appendix C we found reward centering to increase the data efficiency and robustness of Schulman et al.’s (2017) PPO algorithm on continuing versions of several of Todorov et al.’s (2012) Mujoco problems. We expect reward centering would also improve the performance of algorithms that learn no value function at all, such as REINFORCE (Williams, 1992) when applied to continuing problems with eligibility traces, but this has yet to be shown. Many algorithms that were designed for the average-reward criterion already include a form of reward centering, either simple centering (e.g., Tsitsiklis & Van Roy, 1999) or value-based centering (Wan et al., 2021); it is experience with these earlier un-discounted algorithms that led us to explore the utility of reward centering for discounted algorithms. Also expected, but yet to be shown, are the benefits of reward centering with other reinforcement-learning algorithms, including value-based algorithms such as Sarsa (Rummery & Niranjan, 1994), and various offline, actor–critic, and model-based algorithms.
Reward centering is not directly applicable to episodic problems. In these problems the objective is to maximize the sum of rewards only up until the end of an episode; the notion of long-term average reward is undefined and the Laurent-series decomposition (with a state-independent term) no longer holds. Moreover, if reward centering were naively applied to an episodic problem, then it may alter the problem rather than facilitate finding the solution. This is because—unlike in continuing problems—subtracting a constant from all the rewards may change an episodic problem. For example, consider a gridworld where the reward is $-1$ on every step until episode termination upon reaching a goal state. An optimal policy—one that maximizes the total reward per episode—is one that reaches the goal state as soon as possible. However, if the rewards were centered, then the modified rewards would all be zero, and all policies would be equally optimal. The problem would be fundamentally altered by the algorithm that centers (or in general shifts) rewards! The closest thing to reward centering in episodic problems may be the return baseline in policy-gradient methods (e.g., see Sutton & Barto, 2018, Section 13.4), but that may vary from state to state, so the analogy is not really that close. [^2] In continuing problems, reward centering has orthogonal benefits to the baseline or the advantage function in policy-gradient methods. We discuss these and further connections to the literature in Appendix D.
Reward centering may seem similar to, but is different from, a value-function unit with a bias weight (a weight for an input that is always 1). First, the bias weight converges asymptotically to a value that depends on all the other inputs to the value-function unit and that is not in general $r(\pi)/(1-\gamma)$. Second, because the learning of the bias weight interacts with learning all the other weights, we will not obtain the same data efficiency advantages as with reward centering. Reward centering is more akin to the specially learned bias weight in Sutton’s (1988b) NADALINE linear unit. We note that reward centering is also different from but similar in style to methods for adapting the scale of the rewards (e.g., van Hasselt et al., 2016; Pohlen et al., 2018; Schaul et al., 2021), and the two kinds of methods can potentially be used together.
The effects of reward centering on the variance of value estimates is complex. On one hand, reward centering can increase variance because the average-reward estimate changes over time. The simple centering method is particularly susceptible to this in the off-policy setting (e.g., see Figure 3). On the other hand, value-based centering in particular can reduce the variance caused by state-dependent reward changes (cf. Sutton & Barto, 2018, Exercise 10.8). In all cases, optimization techniques could be used to efficiently adapt the step-size parameter of the average-reward estimate (Degris et al., 2024).
@@
Perhaps the most exciting direction for extending reward centering is into new reinforcement learning algorithms that adapt their discount-rate parameter over time. Without reward centering, this would incur huge costs in learning time as the discounted values change by large amounts even for small changes in $\gamma\approx 1$. Most of these changes are due to the state-independent term $r(\pi)/(1-\gamma)$ which, with reward centering, can be adapted instantly to the new value of $\gamma$ using the existing estimate of $r(\pi)$. Concretely, consider the agent has estimated the average reward $\bar{R}$ and the centered discounted value function $\tilde{v}^{\gamma}$ to some level of accuracy. With just this information, the agent can form an estimate of the standard discounted value function corresponding to another discount factor $\gamma^{\prime}$ via $\bar{R}/(1-\gamma^{\prime})+\tilde{v}^{\gamma}$. This is an estimate, of course, but it can be improved quickly with a few samples of experience. In contrast, with standard methods, it would take comparatively longer to raise the estimates to the new mean value and adapt the relative values. Hence, with reward centering, we can imagine efficient methods that adapt their discount factors over time: a low discount rate to learn quickly amidst a lot of uncertainty (like in the beginning of training, or when something about the world changes); a higher discount rate when the world is more predictable to estimate the policy that maximizes the total reward obtained by the agent.
@@ Acknowledgments
The authors gratefully acknowledge funding from NSERC, DeepMind, and the pan-Canadian AI program administered by CIFAR. We thank Huizhen Yu, Arsalan Sharifnassab, Khurram Javed, and other members of the RLAI lab for discussions that helped improve the quality and clarity of the paper. We are also grateful for the computing resources provided by the Digital Research Alliance of Canada.
@@ 14      ${\mathbf{x}}={\mathbf{x}}^{\prime}$
We recommend two small but useful optimizations to these general pseudocodes in Appendix C. The python code is available at github.com/abhisheknaik96/continuing-rl-exps.
@@ Appendix D Connections to Related Approaches
Reward centering and the advantage function have orthogonal benefits. The advantage function benefits the actor by reducing the variance of the updates in the policy space (Sutton & Barto, 2018; Schulman et al., 2016). On the other hand, reward centering benefits the critic’s or baseline’s estimation by eliminating the need to estimate the large state-independent constant offset. Both the quantities involved in the advantage function—${a}_{\pi}^{\gamma}(s,a)={q}_{\pi}^{\gamma}(s,a)-{v}_{\pi}^{\gamma}(s),\forall s,a$—have the large state-independent offset $r(\pi)/(1-\gamma)$. The net effect of the offset is zero when they are subtracted. But the key point is that both the state- and action-value estimates include the large offset. Reward centering removes the need to estimate the large offset for both the state- and action-value function, which simplifies the critic-estimation problem. The actor update is left unchanged with reward centering because the advantage function itself remains unchanged: $\tilde{a}_{\pi}^{\gamma}(s,a)=\tilde{q}_{\pi}^{\gamma}(s,a)-\tilde{v}_{\pi}^{\gamma}(s)$, because $\tilde{q}_{\pi}^{\gamma}(s,a)=q_{\pi}^{\gamma}(s,a)-r(\pi)/(1-\gamma)$ and $\tilde{v}_{\pi}^{\gamma}(s)=v_{\pi}^{\gamma}(s)-r(\pi)/(1-\gamma)$. Hence, we expect reward centering to benefit all the algorithms that estimate values, which include all actor-critic methods that involve advantage estimation.
@@
Reward centering can be seen as reward shaping (Ng et al., 1999) with a constant state-independent potential function: $\Phi(s)=r(\pi)/(1-\gamma),\forall s$. Their Theorem 1 then reiterates that reward centering does not change the optimal policy of the problem. A possible drawback of reward shaping is that fully specifying the potential-based shaping function can be tricky, especially for problems with large state spaces. In the case of reward centering this is relatively easy: the potential function is constant across the entire state space, and we know how to learn the average reward reliably from data.
Finally, we note that the idea of shifting rewards has been explored in episodic problems. Sun et al.’s (2022) experiments show that subtracting a suitable constant from all the rewards can help in some episodic problems. However, we do not expect shifting or centering to help in general in episodic problems: shifting all the rewards by a constant does not change a continuing problem, but can change episodic problems (e.g., the gridworld example from the final section of the main text).